OpenStudy (angelab97):

please help me simplify this x/y-y/x _________ 1/x^2-1/y^2

OpenStudy (angelab97):

preferably by using a picture

OpenStudy (anonymous):

Let's start with the first equation. \[\frac{x}{y}-\frac{y}{x}\] Cross multiply the values. \[\frac{x\times x-y\times y}{yx}\] Simplify the exponents further. Once you've done that, use the difference of squares to simplify again.

OpenStudy (anonymous):

The second equation is \[\frac{1}{{x}^{2}}-\frac{1}{{y}^{2}}\] Again, cross multiply the values. Once you've cross multiplied, use the difference of squares.

OpenStudy (angelab97):

can you please further explain?